# Sideward contact tracing in an epidemic model with mixing groups

### v1

@@ Abstract
We consider a stochastic epidemic model with sideward contact tracing. We assume that infection is driven by interactions within mixing events (gatherings of two or more individuals). Once an infective is diagnosed, each individual who was infected at the same event as the diagnosed individual is contact traced with some given probability. Assuming few initial infectives in a large population, the early phase of the epidemic is approximated by a branching process with sibling dependencies. To address the challenges given by the dependencies, we consider sibling groups (individuals who become infected at the same event) as macro-individuals and define a macro-branching process. This allows us to derive an expression for the effective macro-reproduction number which corresponds to the effective individual reproduction number and represents a threshold for the behaviour of the epidemic. Through numerical illustrations, we show how the reproduction number varies with the mean size of mixing events, the rate of diagnosis and the tracing probability.
@@ 1 Introduction
Contact tracing is recognized as a crucial prevention for pandemic control. Traditional contact tracing involves backward tracing to identify who may have infected the index case and forward tracing to find out who the index case may have infected (individuals who are diagnosed are called index cases in the contact tracing). This paper concerns an innovative tracing method: “sideward” contact tracing. This method is particularly relevant in large gatherings where “superspreading events” (many people get infected at once [lewis2021superspreading]) usually occur. By detecting the gathering, sideward contact tracing can identify individuals infected at the same event as the index case.
@@
The effect of such type of contact tracing has been analyzed, within the framework of simplicial temporal networks, in [mancastroppa2022sideward], where also the traditional forward and backward tracing are considered. In this paper, we focus on the concept of sideward tracing, and we consider the stochastic epidemic model with mixing groups introduced in [ball2022epidemic]. In most epidemic models, infections are assumed to occur between pairs of individuals (one is infectious, and the other is susceptible). Instead of having pairwise interactions, in this model, individuals make contacts via attending mixing events with at least two individuals involved. This allows for the possibility that an infective can infect more than one susceptible on a single occasion and thus provides a suitable framework for the study of sideward tracing.
@@
The mathematical modelling of contact tracing presents several challenges (see an overview in [muller2021contact]). This paper focuses on analyzing the initial behaviour of the epidemic with sideward tracing via a continuous-time branching process incorporating sibling dependencies. Generally speaking, the dependencies introduced by contact tracing complicates the analysis of related branching processes. For instance, in previous studies of models with forward and/or backward contact tracing, lifetimes of siblings are co-dependent on their parents in [ball2011threshold], lifetimes of siblings and parents depend on each other in [zhang2022analysing]. In this paper, the challenge arises from the lifetimes of individuals who were infected in the same mixing event being dependent on each other due to sideward contact tracing.
@@
To address this challenge, we consider the groups of siblings, specifically those who were born (infected) in the same birth (mixing) event, as macro-individuals, such macro-individual behaving independently of each other. The main idea, introduced by [olofsson1997branching, broberg1987sibling], is to consider these macro-individuals as a branching process, referred to as the “macro process”, embedded into the sibling-dependent process. The key point is that macro-individuals reproduce independently, unlike single individuals, as all the dependencies are within the sibling groups. However, our definition of sibling groups differs: the sibling group in [olofsson1997branching] consists of all the children with the same parent, whereas our sibling group comprises children born during the same birth event, so an individual can give birth to multiple sibling groups.
@@
The paper is structured as follows. In Section 2, the epidemic model with mixing groups is defined, as in [ball2022epidemic], and a sideward contact tracing mechanism is introduced in the epidemic model. The approximation of the early stage of the epidemic, with and without contact tracing, is presented in Section 3. In Section 4, the limiting process for the early epidemic with contact tracing is described as a macro branching process, which allows the derivation of macro and individual reproduction numbers in Section 5. A numerical illustration of the effect of sideward tracing on the reproduction number is provided in Section 6. Finally, Section 7 presents the conclusion and discussion.
@@ 2.1 The epidemic model with mixing groups
We describe the SIR epidemic model recently introduced and analysed in [ball2022epidemic, ball2023strong, cortez2024]. Initially, there are $m_{n}$ number of infectives in an otherwise susceptible population of fixed size $n$. Differently from the standard homogeneous SIR epidemic where infections occur through pairwise interactions, infectious contacts in this model are made during temporary gatherings, referred to as “mixing events”, which occur at rate $n\beta$. Each mixing event involves a certain number, distributed as $C^{(n)}$ and independent of other mixing events, of individuals chosen uniformly at random from the population. The mean size of mixing events is denoted by $\mu^{(n)}_{C}=\mathbb{E}\left[C^{(n)}\right]$. Naturally, it is assumed that a mixing event involves at least two individuals and at most $n$ individuals.
Given that a mixing event is of size $c$, i.e. it involves $c$ individuals, any given individual in the population attends the event with probability $c/n$ and any given infectious individual attending the event has probability $\pi_{c}$ of making an infectious contact with any given susceptible individual attending the event, where all such contacts occur independently. Any susceptible individual who is contacted by at least one infective during the event becomes infected and remains infectious for a random period of time $T_{I}\sim Exp(\gamma)$ (with mean $\mathbb{E}\left[T_{I}\right]=1/\gamma$). Further, it is assumed that newly infected individuals cannot transmit the disease within the same event where they were infected, i.e. mixing events are treated as instantaneous, and that there is no latency period. All the processes and random variables described above are assumed to be mutually independent.
@@ 2.2 Sideward contact tracing
This subsection aims to incorporate sideward contact tracing into the epidemic model described above. To this aim, we introduce a diagnosis rate $\delta$, which activates a contact tracing procedure. More precisely, infectives can be removed in three ways: due to natural recovery at rate $\gamma$; due to diagnosis (excluding contact tracing) at rate $\delta$, or through a contact tracing mechanism described in the following. Excluding natural recovery and contact tracing, we use the term “diagnosed” in a broad sense to include various causes of removal that can trigger contact tracing. For example, an individual can be diagnosed due to the onset of symptoms or mass testing.
We assume that when an infectious individual is diagnosed, they are removed (isolated) and the mixing event where they were infected is immediately detected, triggering the sideward contact tracing procedure. Each of the individuals who were infected at the detected event is immediately traced, independently with probability $p$, and they are tested; if infected, they are removed to stop spreading the infection. Table 2.1 lists the important model parameters.
The sideward contact tracing procedure described above only focuses on those infected at the same mixing event as the index case (the one whose diagnosis triggers the tracing), whereas it neglects those who were already infectious at the mixing event. Thus, neither the infector of the index case nor the individuals infected by the index case (infectees) can be traced by this procedure. Moreover, we assume no delays are associated with the tracing process.
@@ 3 Early epidemic approximation
In this section, first, we explain heuristically how the early stages of the epidemic involving mixing groups can be approximated by a branching process; see [ball2023strong] for a more detailed explanation and rigorous proof. Furthermore, we describe how sideward contact tracing modifies the branching process by modifying the lifespan of individuals and by introducing dependencies between siblings who are born in the same birth event (corresponding to individuals who are infected during the same mixing event). The dependencies lead to challenges in the analysis of the reproduction number, which are highlighted here and addressed in the next section.
@@
Under the above asymptotic assumptions and additional integrability conditions, Theorem 3.1 in [ball2023strong] proves that, as the population size $n\to\infty$, the number of infectives in the early stages of the epidemic with mixing groups, without considering sideward contact tracing, is approximated by a branching process $\mathcal{B}$ which we describe in the following.
@@
Denote by $\tilde{Z}_{i},i=1,\dots,G,$ the number of offspring produced at $i-$th birth event, which are i.i.d. random variables, independent of $G$ and equal in distribution to $\tilde{Z}$, described in the following. Given that the size of a mixing group is equal to $c$ which happens with probability $p_{\tilde{C}}(c)$, there are $c-1$ susceptibles in the group, each infected with probability $\pi_{c}$ independently. Thus, the number of individuals infected at an event of size $c$, the number of offspring in a birth event of size $c$, follows a binomial distribution, that is,
$\tilde{Z}|\tilde{C}=c\sim Bin(c-1,\pi_{c}).$
@@
using Equation (2) and (3). A major outbreak in the epidemic is associated to the non-extinction of the approximating branching process $\mathcal{B}$ and occurs with non-zero probability if and only if $R_{0}>1$ [ball2023strong].
@@
$R_{0}=\frac{\beta\mu_{C}}{\gamma+\delta}\mathbb{E}\left[(\tilde{C}-1)\pi_{\tilde{C}}\right]=\frac{\beta\mu_{C}}{\gamma+\delta}\mathbb{E}\left[C(C-1)\pi_{C}\right].$ (5)
@@
Consider one typical individual in $\mathcal{B}_{CT}$ and their siblings, the individuals who were born in the same birth event. Note that individuals who were born from the same parent in other birth events are not called siblings here. The individual dies due to either of the following events happening: natural death (natural recovery in the epidemic) at rate $\gamma$; removal by diagnosis at rate $\delta$; or removal by contact tracing (one of their siblings is removed due to diagnosis and tracing is successful) which happens at rate $(i-1)\delta p$, given that $(i-1)$ of their siblings are currently alive.
@@ 4 The macro branching process
In the limiting process $\mathcal{B}_{CT}$, offspring born at the same birth event constitute a sibling group. While dependency is within each sibling group, sibling groups are independent of their parents and, since each group is produced at independent birth events, sibling groups are independent of each other. By identifying sibling groups as “macro-individual”, we define the macro process $\mathcal{M}$ which is a branching process in terms of independent (macro) individuals.
@@
Each individual in a sibling group gives birth at $\beta\mu_{C}$, thus, at age $t$ the whole group gives birth to a new sibling group at a total rate $Y(t)\beta\mu_{C}$. The whole sibling group dies when the Markov process $\{Y(t)\}_{t\geq 0}$ reaches state 0. We can thus simply describe the macro branching process $\mathcal{M}$ as a Crump-Mode-Jagers branching process where the birth rate of each individual at age $t$ is independent on other individuals and distributed as $\lambda(t):=Y(t)\beta\mu_{C}$. See [jagers1975] for an overview and details on Crump-Mode-Jagers branching processes.
@@
Remark 4.1.
For the macro process $\mathcal{M}$, it is also possible to easily express the generation time distribution by normalising $\mathbb{E}\left[\lambda(t)\right]$ so that it integrates to $1$ (dividing by the reproduction number) and to express the Malthusian parameter, or growth rate, $r_{\mathcal{M}}$ as the unique solution of
[…]
This provides additional tools that can be used for further analyses, left to future work. For example, it is known [favero2022] that some preventive measures (as isolation, mass testing, forward and backward contact tracing) affect the generation time distribution leading to biased estimates of the reproduction number. This suggests that a similar study on the effect of sideward contact tracing on the generation time distribution in the epidemic model and resulting biases might yield interesting results.
Remark 4.2.
where $\sigma_{i}(s)$ is the time of birth of the $i^{th}$ macro-individual alive at time $s$, with $s-\sigma_{i}(s)$ thus being its age. By interpreting the size process $Y$ as a “characteristic” of a macro-individual, the expression above allows us to interpret $I_{\mathcal{B}_{CT}}(s)$ as the “total characteristic” of $\mathcal{M}$ at time $s$, as in [jagers1984]. Then, the classical theory of Crump-Mode-Jagers branching processes [jagers1984] implies in the supercritical case that, for large times, $I_{\mathcal{M}}(s)\approx e^{r_{\mathcal{M}}s}W$, with $W$ being a positive random variable; and that $I_{\mathcal{B}_{CT}}(s)\approx e^{r_{\mathcal{M}}s}m_{Y}W$, with $m_{Y}$ being a positive constant.
@@ 5.1 Effective macro reproduction number
with $m_{i0}$ in Equation (13) and $Y_{0}\sim Bin(\tilde{C}-1,\pi_{\tilde{C}})$. We have thus proved the following Theorem.
@@
$R_{e}=\frac{\beta}{\gamma+\delta}\sum_{c=2}^{\infty}c\mathbb{P}(C=c)\sum_{k=1}^{c-1}m_{i0}\binom{c-1}{i}\pi_{c}^{i}(1-\pi_{c})^{c-i-1},$ (14)
@@ 5.3 Extinction probability
In the following, we obtain an expression for the extinction probability of the branching process $\mathcal{B}_{CT}$, denoted by $z_{\mathcal{B}_{CT}}$ when there is one ancestor, and equal to $z_{\mathcal{B}_{CT}}^{m}$ when there are $m$ ancestors. Consider now the process $\mathcal{M}$ initiated with one macro-individual and denote by $z_{\mathcal{M}}$ its extinction probability. By the standard theory of branching processes, $z_{\mathcal{M}}$ is the smallest solution in $[0,1]$ of $f_{H}(s)=s$, with $f_{H}(s)$ being the offspring probability generating function for $\mathcal{M}$. More explicitly,
@@
To obtain the extinction probability $z_{\mathcal{B}_{CT}}$, we condition on the number of birth events (sibling groups/macro-individuals) produced by the single ancestor in $\mathcal{B}_{CT}$ being equal to $g$. Since the ancestor cannot be contact traced, but can recover or be diagnosed, this number is geometrically distributed as $G$ in (1) with success probability $\frac{\gamma+\delta}{\gamma+\delta+\beta\mu_{C}}$. We then consider the macro process $\mathcal{M}$ initiated with $g$ macro-individuals, which has extinction probability $z_{\mathcal{M}}^{g}$. Therefore, the extinction probability of the process $\mathcal{B}_{CT}$ starting with one ancestor is given by
$z_{\mathcal{B}_{CT}}=\mathbb{E}\left[z_{\mathcal{M}}^{G}\right]=\frac{\gamma+\delta}{\gamma+\delta+\beta\mu_{C}(1-z_{\mathcal{M}})},$ (20)
and hence in the epidemic setting with $m$ initial infectives, a major outbreak occurs with probability $1-z_{\mathcal{B}_{CT}}^{m}$.
@@
6 Numerical illustration
The goal of this section is to illustrate the effect of sideward contact tracing by means of some numerical examples relying on the theoretical results of the previous section.

### v2

@@ Abstract
We consider a stochastic epidemic model with sideward contact tracing. We assume that infection is driven by interactions within mixing events (gatherings of two or more individuals). Once an infective is diagnosed, each individual who was infected at the same event as the diagnosed individual is contact traced with some given probability. Assuming few initial infectives in a large population, the early phase of the epidemic is approximated by a branching process with sibling dependencies. To address the challenges given by the dependencies, we consider sibling groups (individuals who become infected at the same event) as macro-individuals and define a macro-branching process. This allows us to derive an expression for the effective macro-reproduction number which corresponds to the effective individual reproduction number and represents a threshold for the behaviour of the epidemic. Through numerical examples, we show how the reproduction number varies with the distribution of the mixing event size, the mean size, the rate of diagnosis and the tracing probability.
@@ 1 Introduction
Contact tracing is recognized as a crucial prevention for pandemic control. Traditional contact tracing involves backward tracing to identify who may have infected the index case and forward tracing to find out who the index case may have infected (individuals who are diagnosed are called index cases in the contact tracing). This paper concerns an innovative tracing method known as “sideward” contact tracing, first introduced and analysed in [15] within the framework of simplicial activity-driven temporal networks. We focus on a related sideward tracing strategy which identifies individuals attending the same gathering as the index case. This mechanism is particularly relevant in the context of large gatherings where “superspreading events” (many people get infected at once [14]) usually occur.
@@
While [15] makes a significant step in the analysis of the effects of sideward tracing by focusing on an epidemic model spreading on a temporal network, we instead focus on a different framework arising from classical stochastic SIR models (see, e.g., [5] for an overview). In particular, we adopt the stochastic epidemic model with mixing groups introduced in [1]. In most epidemic models, infections are assumed to occur between pairs of individuals (one is infectious, and the other is susceptible). In this model, instead of having pairwise interactions, individuals make contacts via attending mixing events with at least two individuals involved. This allows for the possibility that an infective can infect more than one susceptible on a single occasion and thus provides a suitable framework for the study of sideward tracing.
The mathematical modelling of contact tracing presents several challenges (see an overview in [17]). This paper focuses on analysing the initial behaviour of the epidemic with sideward tracing via a continuous-time branching process incorporating sibling dependencies. In this context, an individual in the branching process corresponds to an infectious individual in the epidemic setting, while their offspring represent the secondary cases they infect and siblings are those who are infected by the same individual. Notably, sibling dependencies are not limited to epidemic settings. They are also highly relevant in biological population models, where they play a key role in understanding population dynamics [18]. A common example is sibling competition, which arises when siblings share limited resources [19], leading to a negative correlation in the number of offspring each sibling can successfully produce.
@@
Generally speaking, the dependencies introduced by contact tracing complicate the analysis of related branching processes. For instance, in previous studies of models with forward and/or backward contact tracing, lifetimes of siblings are co-dependent on their parents in [3], lifetimes of siblings and parents depend on each other in [20]. In this paper, the challenge arises from the lifetimes of individuals who were infected in the same mixing event being dependent on each other due to sideward contact tracing.
@@
To address this challenge, we consider the groups of siblings, specifically those who were born (infected) in the same birth (mixing) event, as macro-individuals, such macro-individual behaving independently of each other. The main idea, introduced by [18, 6], is to consider these macro-individuals as a branching process, referred to as the “macro process”, embedded into the sibling-dependent process. The key point is that macro-individuals reproduce independently, unlike single individuals, as all the dependencies are within the sibling groups. However, our definition of sibling groups differs: the sibling group in [18] consists of all the children with the same parent, whereas our sibling group comprises children born during the same birth event, so an individual can give birth to multiple sibling groups.
@@
The paper is structured as follows. In Section 2, the epidemic model with mixing groups is defined, as in [1], and a sideward contact tracing mechanism is introduced in the epidemic model. The approximation of the early stage of the epidemic, with and without contact tracing, is presented in Section 3. In Section 4, the limiting process for the early epidemic with contact tracing is described as a macro branching process, which allows the derivation of macro and individual reproduction numbers in Section 5. A numerical illustration of the effect of sideward tracing on the reproduction number is provided in Section 6. Finally, Section 7 presents the conclusion and discussion.
@@ 2.1 The epidemic model with mixing groups
We describe the SIR epidemic model recently introduced and analysed in [1, 2, 7]. Initially, there are $m_{n}$ number of infectives in an otherwise susceptible population of fixed size $n$. Differently from the standard homogeneous SIR epidemic where infections occur through pairwise interactions, infectious contacts in this model are made during temporary gatherings, referred to as “mixing events”, which occur at rate $n\beta$. Each mixing event involves a certain number, distributed as $C^{(n)}$ and independent of other mixing events, of individuals chosen uniformly at random from the population. The mean size of mixing events is denoted by $\mu^{(n)}_{C}=\mathbb{E}\left[C^{(n)}\right]$. Naturally, it is assumed that a mixing event involves at least two individuals and at most $n$ individuals.
Given that a mixing event is of size $c$, i.e. it involves $c$ individuals, any given individual in the population attends the event with probability $c/n$ and any given infectious individual attending the event has probability $\pi_{c}$ of making an infectious contact with any given susceptible individual attending the event, where all such contacts occur independently. The dependence of $\pi_{c}$ on $c$ is designed to capture how event size influences the likelihood of transmission in real-world scenarios. For instance, individuals in larger gatherings may adopt more cautious behaviours, such as maintaining physical distance, which can lower the transmission risk. Conversely, smaller events tend to foster closer interactions, such as prolonged conversations, which can be more conducive to transmission.
Any susceptible individual who is contacted by at least one infective during the event becomes infected and remains infectious for a random period of time $T_{I}\sim Exp(\gamma)$ (with mean $\mathbb{E}\left[T_{I}\right]=1/\gamma$). Further, it is assumed that newly infected individuals cannot transmit the disease within the same event where they were infected, i.e. mixing events are treated as instantaneous, and that there is no latency period. All the processes and random variables described above are assumed to be mutually independent.
@@ 2.2 Sideward contact tracing
This subsection aims to incorporate sideward contact tracing into the epidemic model described above. To this aim, we introduce a diagnosis rate $\delta$, which activates a contact tracing procedure. More precisely, infectives can be removed in three ways: due to natural recovery at rate $\gamma$; due to diagnosis (excluding contact tracing) at rate $\delta$, or through a contact tracing mechanism described in the following. For clarity, the term “diagnosed” is used broadly here to include various causes of removal that can initiate contact tracing. For example, an individual can be diagnosed due to the onset of symptoms or mass testing.
We assume that when an infectious individual is diagnosed, they are removed (isolated) and the mixing event where they were infected is immediately detected, triggering the sideward contact tracing procedure. Such individual is referred to as the index case. Ideally, contact tracing would identify all of the events involving the index case and individuals present at the detected events, including those who infected the index case (backward tracing) and those they subsequently infected (forward tracing). However, such a comprehensive hybrid approach introduces additional dependencies between infectors and infectees, making the analysis significantly more challenging. In this paper, we focus exclusively on the novel mechanism of sideward tracing, not only because of its mathematical tractability, but also to better analyse its effect separately from the effect of the better-known mechanisms of backward and forward tracing. Specifically, we assume that sideward tracing identifies only the event where the index case was infected and the individuals present at that event, while excluding both the infector(s) and infectees.
This assumption in principle includes both individuals infected at the same event as the index case and those infected at distinct events but present at the detected event. However, in the initial phase of the epidemic, the latter type of individuals is not present because mixing events that involve at least one infective include, almost surely, only one infective with all others being susceptible (see Section 3). Our analysis concerns the initial phase of the epidemic, consequently, we can safely simplify the sideward tracing mechanism to tracing only those infected at the same event as the index case, while neglecting those present but infected elsewhere.
For mathematical formulation, we define the sideward tracing procedure as follows: each individual who is infected at the same event as the index case is traced, independently with probability $p$, and tested. If found infectious, they are immediately removed to stop spreading the infection. Table 2.1 lists the important model parameters.
@@ 3 Early epidemic approximation
In this section, first, we explain heuristically how the early stages of the epidemic involving mixing groups can be approximated by a branching process; see [2] for a more detailed explanation and rigorous proof. Furthermore, we describe how sideward contact tracing modifies the branching process by modifying the lifespan of individuals and by introducing dependencies between siblings who are born in the same birth event (corresponding to individuals who are infected during the same mixing event). The dependencies lead to challenges in the analysis of the reproduction number, which are highlighted here and addressed in the next section.
@@
Under the above asymptotic assumptions and additional integrability conditions, Theorem 3.1 in [2] proves that, as the population size $n\to\infty$, the number of infectives in the early stages of the epidemic with mixing groups, without considering sideward contact tracing, is approximated by a branching process $\mathcal{B}$ which we describe in the following.
@@
Denote by $\tilde{Z}_{i},i=1,\dots,G,$ the number of offspring produced at $i-$th birth event, which are i.i.d. random variables, independent of $G$ and equal in distribution to $\tilde{Z}$, described in the following. Given that the size of a mixing group is equal to $c$ which happens with probability $p_{\tilde{C}}(c)$, there are $c-1$ susceptibles in the group, each infected with probability $\pi_{c}$ independently. Thus, the number of individuals infected at an event of size $c$, the number of offspring $\tilde{Z}|\tilde{C}=c$ in a birth event of size $c$, follows a binomial distribution $Bin(c-1,\pi_{c})$. Consequently, $\tilde{Z}$ follows a mixed-binomial distribution, that is,
$\tilde{Z}\sim MixBin(\tilde{C}-1,\pi_{\tilde{C}}).$
@@
using Equation (2) and (3). A major outbreak in the epidemic is associated to the non-extinction of the approximating branching process $\mathcal{B}$ and occurs with non-zero probability if and only if $R_{0}>1$ [2].
@@
$R_{0}=\frac{\beta\mu_{C}}{\gamma+\delta}\mathbb{E}\left[(\tilde{C}-1)\pi_{\tilde{C}}\right]=\frac{\beta}{\gamma+\delta}\mathbb{E}\left[C(C-1)\pi_{C}\right].$ (5)
@@
Consider one typical individual in $\mathcal{B}_{CT}$ and their siblings, the individuals who were born in the same birth event. Note that individuals who were born from the same parent in other birth events are not called siblings here. The individual dies due to either of the following events happening: natural death (natural recovery in the epidemic) at rate $\gamma$; removal by diagnosis at rate $\delta$; or given that $(i-1)$ of their siblings are currently alive, one of their siblings is removed due to diagnosis and tracing is successful which happens at rate $(i-1)\delta p$.
@@ 4 The macro branching process
In the limiting process $\mathcal{B}_{CT}$, offspring born at the same birth event constitute a sibling group. While dependency is within each sibling group, sibling groups are independent of their parents and, since each group is produced at independent birth events, sibling groups are independent of each other. By identifying sibling groups as “macro-individual”, we define the macro process $\mathcal{M}$ which is a branching process in terms of independent (macro) individuals. Figure 4.1 shows an example how the individual process and macro process are linked.
Figure 4.1: Illustration of an individual process $\mathcal{B}_{CT}$ and its corresponding macro process $\mathcal{M}$. On the left, circles represent individuals grouped into sibling groups, enclosed within dashed rectangles. The macro process on the right simplifies the individual process by aggregating sibling groups into macro-individuals, represented as squares labeled by the respective sibling group sizes.
@@
As a consequence, from a non-zero state $Y(t)=i$, two types of jumps can occur. The first two events both result in the same type of jump: the group size is decreased by $1$ ($i\to i-1$), whereas the third one leads to another type of jump where the group size is decreased by more than 1 ($i\to j$). In conclusion, the transition rates of the process from state $i$ to $j$, denoted by $q_{i,j}$, are as follows:
@@
Each individual in a sibling group gives birth at $\beta\mu_{C}$, thus, at age $t$ the whole group gives birth to a new sibling group at a total rate $Y(t)\beta\mu_{C}$. The whole sibling group dies when the Markov process $\{Y(t)\}_{t\geq 0}$ reaches state 0. We can thus simply describe the macro branching process $\mathcal{M}$ as a Crump-Mode-Jagers branching process where the birth rate of each individual at age $t$ is independent on other individuals and distributed as $\lambda(t):=Y(t)\beta\mu_{C}$. See [9] for an overview and details on Crump-Mode-Jagers branching processes.
@@ 5.1 Effective macro reproduction number
with $m_{i0}$ in Equation (15) and $Y_{0}\sim MixBin(\tilde{C}-1,\pi_{\tilde{C}})$. We have thus proved the following Theorem.
@@
$R_{e}=\frac{\beta}{\gamma+\delta}\sum_{c=2}^{\infty}c\mathbb{P}(C=c)\sum_{i=1}^{c-1}m_{i0}\binom{c-1}{i}\pi_{c}^{i}(1-\pi_{c})^{c-i-1},$ (16)
@@ 5.3 Extinction probability
In the following, we obtain an expression for the extinction probability of the branching process $\mathcal{B}_{CT}$, denoted by $z_{\mathcal{B}_{CT}}$ when there is one ancestor, and equal to $z_{\mathcal{B}_{CT}}^{m}$ when there are $m$ ancestors. Consider now the process $\mathcal{M}$ initiated with one macro-individual and denote by $z_{\mathcal{M}}$ its extinction probability. By the standard theory of branching processes, $z_{\mathcal{M}}$ is the smallest solution in $[0,1]$ of $f_{H}(s)=s$, with $f_{H}(s)$ being the offspring probability generating function for $\mathcal{M}$ and $H$ defined in Equation (11). More explicitly,
@@
To obtain the extinction probability $z_{\mathcal{B}_{CT}}$, we condition on the number of birth events (sibling groups/macro-individuals), denoted by $J$, produced by the single ancestor in $\mathcal{B}_{CT}$ being equal to $j$. We then consider the macro process $\mathcal{M}$ initiated with $j$ macro-individuals, which has extinction probability $z_{\mathcal{M}}^{j}$. This implies that
$z_{\mathcal{B}_{CT}}=\mathbb{E}\left[z_{\mathcal{M}}^{J}\right]=\sum_{j=0}^{\infty}\mathbb{E}\left[z_{\mathcal{M}}^{J}|J=j\right]\mathbb{P}(J=j)=\sum_{j=0}^{\infty}z_{\mathcal{M}}^{j}\mathbb{P}(J=j)$ (22)
Since the ancestor cannot be contact traced, but can recover or be diagnosed, $J$ is geometrically distributed with success probability $(\gamma+\delta)/(\gamma+\delta+\beta\mu_{C})$ similar to distribution of $G$ in Equation (1). That is, we have
$\mathbb{P}(J=j)=\left(\frac{\beta\mu_{C}}{\gamma+\delta+\beta\mu_{C}}\right)^{j}\frac{\gamma+\delta}{\gamma+\delta+\beta\mu_{C}},\qquad(j=0,1,...).$
Embedding the equation above into Equation (22) gives
$z_{\mathcal{B}_{CT}}=\frac{\gamma+\delta}{\gamma+\delta+\beta\mu_{C}}\sum_{j=0}^{\infty}\left(\frac{z_{\mathcal{M}}\beta\mu_{C}}{\gamma+\delta+\beta\mu_{C}}\right)^{j}=\frac{\gamma+\delta}{\gamma+\delta+\beta\mu_{C}(1-z_{\mathcal{M}})}.$ (23)
Note that in the epidemic setting with $m$ initial infectives, a major outbreak occurs with probability $1-z_{\mathcal{B}_{CT}}^{m}$.
@@
6 Numerical results
The goal of this section is to illustrate the effect of sideward contact tracing by means of some numerical examples relying on the theoretical results of the previous section. In particular, we consider two forms of infection probability: one being fixed, i.e., $\pi_{c}\equiv\pi$ and another depends on the size, defined as $\pi_{c}=\frac{2}{c}$ for all $c\geq 2$. We consider this form for two main reasons. First, in an event involving only two people, the probability of one infectious individual transmitting the infection to the other is effectively 1, as in the standard stochastic SIR model. Second, this form serves as an example to illustrate how infection probability decreases as the event size increases. In larger events, an infective is likely to interact with more people, potentially dispersing the risk of close contact with any single individual. Notably, when $\pi_{c}={2}/{c}$, the basic reproduction number with diagnosis $R_{0}$ given by Equation (5) equals ${2\beta(\mu_{C}-1)}/{(\gamma+\delta)}.$
Additionally, we consider three different distribution of $C$ all chosen to have the same mean $\mu_{C}$ (following the choice in [1]):
  1. (i)
    $C$ follows a logarithmic distribution with
    $\mathbb{P}(C=c)=\frac{(1-\alpha)^{c}}{(-\log(\alpha)-(1-\alpha))c},\quad(c=2,3,...)$
    and in this case, the mean event size $\mu_{C}=\frac{(1-\alpha)^{2}}{(-\log(\alpha)-(1-\alpha))\alpha}$.
  2. (ii)